# String Decomposition and Gravitational Waves in High-quality Axion Gauge Theories

### v1

@@ Abstract
Remarkably, Eq. (10) implies that any string with axion winding number $|w_{a}|>1$ is topologically equivalent to $w_{a}$ unit-winding axion strings plus $c$ pure gauge strings. To deliver this point explicitly, we survey a number of concrete decomposition examples in table. 1. In all cases, at least one pure gauge string is needed for the decomposition to work, i.e., $c\neq 0$. This property continues to holds for most strings with higher $(m,n)$ values not shown in the table. We choose $p+q\geq 10$ such that Eq. (3) allows for a high quality axion, after taking into account of astrophysical constraints $f_{S},f_{T}>f_{a}>10^{9}\,$GeV Adams et al. (2022).
@@
In the above discussion of depleting high $|w_{a}|$ strings to avoid the domain wall problem, a key player is the need for pure gauge strings. There are two epochs when the gauge strings prevail throughout the universe. One happens early on $t_{S}<t<t_{T}$, where only the $S$ field develops a VEV thus only gauge strings are present. The second epoch corresponds to $t>t_{\rm QCD}$ scale where the axion domain walls decompose the strings following the rule of Eq. (10). The decomposition liberates gauge strings. While all the resulting axion strings are quickly destroyed, gauge strings do not talk to axion domain walls and survive until later times. On the other hand, during $t_{T}<t<t_{\rm QCD}$, almost all strings filling the universe are axion strings with $w_{a}\neq 0$ (see e.g. table. 1).
@@
$W(t_{i})=\Theta\left(t_{i}-t_{S}\right)\Theta\left(t_{T}-t_{i}\right)+\Theta\left(t_{i}-t_{\rm QCD}\right)\ .$ (15)
Throughout the calculation, we assume the universe is radiation dominated at early times. At time $t_{S,T}$, the universe has a temperature $f_{S,T}$, respectively.
The resulting GW spectrum is shown in Fig. 2. It features two plateaus and a valley in between, which can be understood as the following. Using Eq. (13), one can show that gauge string loops produced around a time $t_{i}$ make a triangle shaped GW spectrum in the log-log plot of $\Omega_{\rm GW}$ versus $f$, peaked at $f_{*}\simeq 10^{13}\,{\rm Hz}\times[T(t_{i})/f_{S}]$, where $T(t_{i})$ denotes the temperature of the universe at time $t_{i}$. It corresponds to the two terms in the square bracket of Eq. (14) are comparable. In the $\alpha\gg\Gamma Gf_{S}^{2}$ limit, GW is dominantly produced with original frequency $(\alpha t_{i})^{-1}$ around time $t\simeq\alpha t_{i}/(\Gamma Gf_{S}^{2})$. Thus a larger $f_{S}$ brings $t$ closer to $t_{i}$, resulting in more redshift till today and smaller $f_{*}$. The spectrum grows as $f^{3}$ for $f<f_{*}$ and falls as $1/f$ for $f>f_{*}$. For string loop production lasting over an extended period, the corresponding triangles stack next to each other to form a plateau. The two stages of gauge string production discussed above leads to two plateaus. With $T(t_{\rm QCD})\simeq 0.1\,$GeV, the lower-$f$ plateau extends up to a frequency around $f\simeq 1\,{\rm Hz}\times(10^{12}\,{\rm GeV}/f_{S})$. The higher-$f$ plateau spans a window $10^{13}\,{\rm Hz}\times(f_{T}/f_{S})\lesssim f\lesssim 10^{13}\,{\rm Hz}$. For frequencies in between, the GW spectrum is suppressed due to the lack of gauge strings during $t_{T}<t_{i}<t_{\rm QCD}$. With the astrophysical bound $f_{T}\sim f_{a}>10^{9}\,$GeV, the valley always exists. Therefore, the shape of $\Omega_{\rm GW}h^{2}$ shown in Fig. 2 resembles a robust prediction of the Barr-Seckel models with any coprime $(p,q)$ pairs.

### v2

@@ Abstract
Remarkably, Eq. (10) implies that any string with axion winding number $|w_{a}|>1$ is topologically equivalent to $w_{a}$ unit-winding axion strings plus $c$ pure gauge strings. To deliver this point explicitly, we survey a number of concrete decomposition examples in table. 1. In all cases, at least one pure gauge string is needed for the decomposition to work, i.e., $c\neq 0$. This property continues to hold for most strings with higher $(m,n)$ values not shown in the table. We choose $p+q\geq 10$ such that Eq. (3) allows for a high quality axion, after taking into account of astrophysical constraints $f_{S},f_{T}>f_{a}>10^{9}\,$GeV Adams et al. (2022).
@@
In the above discussion of depleting high $|w_{a}|$ strings to avoid the domain wall problem, a key player is the need for pure gauge strings. There are two epochs when the gauge strings prevail throughout the universe. One happens early on $t_{S}<t<t_{T}$, where only the $S$ field develops a VEV thus only gauge strings are present. The second epoch corresponds to $t>t_{2}$ scale where the axion domain walls decompose the strings following the rule of Eq. (10). The decomposition liberates gauge strings. While all the resulting axion strings are quickly destroyed, gauge strings do not talk to axion domain walls and survive until later times. On the other hand, during $t_{T}<t<t_{2}$, almost all strings filling the universe are axion strings with $w_{a}\neq 0$ (see e.g. table. 1).
@@
$W(t_{i})=\Theta\left(t_{i}-t_{S}\right)\Theta\left(t_{T}-t_{i}\right)+\Theta\left(t_{i}-t_{2}\right)\ ,$ (15)
where $t_{2}$ is the axion string-wall collapsing time and given by Eq. (12). Throughout the calculation, we assume the universe is radiation dominated at early times. At time $t_{S,T}$, the universe has a temperature $f_{S,T}$, respectively.
The resulting GW spectrum is shown in Fig. 2. It features two plateaus and a valley in between, which can be understood as the following. Using Eq. (13), one can show that gauge string loops produced around a time $t_{i}$ make a triangle shaped GW spectrum in the log-log plot of $\Omega_{\rm GW}$ versus $f$, peaked at $f_{*}\simeq 10^{13}\,{\rm Hz}\times[T(t_{i})/f_{S}]$, where $T(t_{i})$ denotes the temperature of the universe at time $t_{i}$. It corresponds to the two terms in the square bracket of Eq. (14) are comparable. In the $\alpha\gg\Gamma Gf_{S}^{2}$ limit, GW is dominantly produced with original frequency $(\alpha t_{i})^{-1}$ around time $t\simeq\alpha t_{i}/(\Gamma Gf_{S}^{2})$. Thus a larger $f_{S}$ brings $t$ closer to $t_{i}$, resulting in more redshift till today and smaller $f_{*}$. The spectrum grows as $f^{3}$ for $f<f_{*}$ and falls as $1/f$ for $f>f_{*}$. For string loop production lasting over an extended period, the corresponding triangles stack next to each other to form a plateau. The two stages of gauge string production discussed above leads to two plateaus. With $T(t_{2})\leq T(t_{\rm QCD})\simeq 0.1\,$GeV, the lower-$f$ plateau extends up to a frequency $f\lesssim 1\,{\rm Hz}\times(10^{12}\,{\rm GeV}/f_{S})$. The higher-$f$ plateau spans a window $10^{13}\,{\rm Hz}\times(f_{T}/f_{S})\lesssim f\lesssim 10^{13}\,{\rm Hz}$. For frequencies in between, the GW spectrum is suppressed due to the lack of gauge strings during $t_{T}<t_{i}<t_{2}$. With the astrophysical bound $f_{T}\sim f_{a}>10^{9}\,$GeV, the valley always exists. Therefore, the shape of $\Omega_{\rm GW}h^{2}$ shown in Fig. 2 resembles a robust prediction of the Barr-Seckel models with any coprime $(p,q)$ pairs.